OpenStudy (anonymous):

i need help factoring thie expression by difference of squares a^2 + 4a + 4 - b^2

OpenStudy (mertsj):

\[a^2-4a+4=(a-2)^2\]

OpenStudy (mertsj):

Does that help?

OpenStudy (anonymous):

ya but how do u type out (a-2)2

OpenStudy (mertsj):

Use the equation button below the box

OpenStudy (anonymous):

i mean what would be the long form of (a-2)^2

OpenStudy (mertsj):

\[[(a-2)^2]-b^2=[(a-2)+b][(a-2)-b]\]

OpenStudy (mertsj):

The expanded form of (a-2)^2 is a^2-4a+4. Remember?? That is what we had before we factored it.

OpenStudy (anonymous):

ook